# Conditioning LLMs to Generate Code-Switched Text

### v2

@@ Abstract
Code-switching (CS) is still a critical challenge in Natural Language Processing (NLP), due to the limited availability of large-scale, diverse CS datasets for robust training and evaluation. Despite recent advances, the true capabilities and limitations of LLMs in handling CS remain underexplored. In this work, we investigate the extent to which LLMs can be used in a framework for CS text generation, focusing on the English-Spanish language pair. Our proposed methodology consists of back-translating natural CS sentences into monolingual English, and using the resulting parallel corpus to fine-tune LLMs to turn monolingual sentences into CS. We thoroughly analyse the models’ performance through a study on human preferences, a qualitative error analysis, an evaluation with popular reference-based metrics and LLM-based judgment. Results show that fine-tuning can be a key step to ensure that current LLMs consistently generate fluent code-switched text and that our methodology generates high-quality outputs, expanding research opportunities in CS communication. We find that traditional metrics do not correlate with human judgement when assessing the quality of the generated CS data but LLM-based judgment aligns more closely with human preferences. We release our code and generated dataset under a CC-BY-NC-SA license. [^1] URL to be announced upon acceptance.
1 Introduction
@@
Despite the prevalence of code-switching, most research in Natural Language Processing (NLP) assumes monolingualism as a standard for human communication. However, this implicit decision means that state-of-the-art models are not able to properly interpret or generate CS data. Even advances in multilingual language modelling (Lin et al., 2022; Chowdhery et al., 2023) have not led to significant improvements, and performance on CS data is still poor compared to performance on monolingual data (Aguilar et al., 2020; Winata et al., 2021). This occurs because there is little CS text available in the multilingual pretraining data. Similarly, there are no parallel datasets available to learn to generate CS in a supervised fashion, as one would expect for tasks such as machine translation. Finally, existing methodology for evaluating automatically generated CS text, which has specific needs different from other text generation tasks, are still not good enough and fail to capture nuances of CS text (Srivastava and Singh, 2021). It is therefore crucial to develop methodologies to enable models to generate natural CS text and simultaneously implement robust evaluation frameworks that can assess how well NLP systems handle CS across multiple tasks. We argue that both of these goals require models that can conditionally generate CS from monolingual text. Consequently, our research focuses on the development of a methodology to fine-tune and evaluate LLMs on the task of CS generation, following three main research questions:
@@
Do automatic metrics for Natural Language Generation (NLG) or LLM judges correlate well with human judgement for the task of CS generation?
Based on these research questions, we propose a novel approach to generate CS from monolingual text using LLMs and apply it to the English-Spanish pair. We create a new parallel English–CS corpus, EN-CS, by leveraging natural CS data and using LLMs to perform back-translation from CS into English, resulting in high-quality pseudo-parallel pairs, suitable for training and evaluating models on CS generation (RQ2). We provide a comprehensive comparison of CS generation using LLMs in both zero-shot and fine-tuned settings, and we compare their performance against that of a dedicated machine translation (MT) model (RQ1). Finally, we evaluate our methodology both qualitatively, with a study on human preferences and a manual error analysis, and quantitatively, using automatic NLG metrics and LLM as a judge, which allows us to study the correlation between human and automatic evaluation for this task (RQ3).
2 Related Work
@@
Most code-switched data stems from social media, while other popular data sources include recordings and transcriptions (Winata et al., 2023). Shared tasks using such CS data have been organized for the tasks of Language Identification (Solorio et al., 2014; Molina et al., 2016) and Sentiment Analysis (Patwa et al., 2020). Similarly, two benchmarks exist to evaluate model performance on CS text, covering different language pairs and tasks: LINCE (Aguilar et al., 2020), which covers tasks such as Part Of Speech tagging or Sentiment Analysis; and GLUECoS (Khanuja et al., 2020), which focuses on NLU tasks for Hindi-English. Unfortunately, GLUECoS cannot be currently used without access to the X API.
@@
CS generation has seldom been tackled in previous research. Approaches include linguistically informed techniques to find plausible switching points (Pratapa et al., 2018; Gupta et al., 2020; Gregorius and Okadome, 2022; Hsu et al., 2023), data augmentation (Tarunesh et al., 2021) and, more recently, prompting LLMs for CS generation (Yong et al., 2023). While CS generation is often evaluated by human annotators (Tarunesh et al., 2021; Gregorius and Okadome, 2022), there remains a need for robust automatic evaluation methodologies to assess the naturalness and fluency of the generated texts, with some recent studies already exploring approaches like Judge-LLMs (Kuwanto et al., 2024).
3 Parallel Data Creation
@@
3.1 The LINCE benchmark
[…]
estaba aquí three feet away .
spa spa eng eng eng eng&spa
[…]
LINCE comprises around $95,000$ train, $20,000$ development, and $33,000$ test instances for the English-Spanish pair. We deduplicate the instances among splits, and filter and pre-process the instances to ensure that they are suitable for our task by removing links, replacing usernames with the placeholder <user>, and detokenizing all instances with the script provided as part of the Moses toolkit (Koehn et al., 2007). After this preprocessing, we obtain a more natural version of the LINCE data. A preliminary analysis reveals that many sentences in LINCE are monolingual or contain a single word in one language that often correspond to a borrowing, as shown in Example 3.1. In order to ensure that all of our sentences actually contain CS, we filter sentences that do not have at least two words in each language.
[…]
After these pre-processing and filtering steps, we end up with $12,933$ train, $2,461$ development and $,35$ 353 test instances. The comparison between the original size of LINCE and the final number of sentences selected for our experiments after pre-processing is shown in Table 1.
Train Dev Test
Original 94,728 19,574 33,361
Pre-processed 12,933 2,461 5,353
EN-CS 10,703 791 1,040
@@ 3.1 The LINCE benchmark
3.2 EN-CS
The next step in our method requires creating a pseudo-parallel English-CS dataset by translating the natural code-switched instances into monolingual text. As there are no available machine translation systems to convert from English-Spanish CS text to English monolingual text, we instead make use of prompt engineering, using the Command R model (Cohere For AI, 2024), one of the strongest publicly available models at the time.
We perform an initial set of experiments to determine the optimal prompt to generate monolingual English versions of the code-switched data. Ideally, we aim for a prompt that generates translations that maintain the meaning of the original sentences, are fluent and natural, whose grammar is correct, and that do not contain any Spanish words or phrases. After extensive testing (see Appendix A), we use the following prompt in a 5-shot setting: Now convert this code-switched phrase to English. Leave the parts in English as they are, focus on translating the parts in Spanish. Finally, we filter output instances that contain profanity that was not present in the source texts or irrelevant information, such as “Of course, here’s your translation:”.
In order to create a valid gold standard test set, we perform a manual post-edition of the the monolingual test translations for $1,040$ instances of the LINCE test set. The post-edition was carried out by three proficient speakers of English and Spanish, who were provided with specific guidelines as shown in Appendix B.
@@ 3.2 EN-CS
3.2.1 Quality assessment
To estimate the quality of the post-edition process, we compare the post-editions of two annotators on 100 additional random instances. The results show that 75% of the sentences remain unchanged, as they are already adequate. There is a 87.87% similarity between the post-editions of the two annotators, as measured by Levenshtein distance, demonstrating a high degree of consistency and quality in the post-edition process.
4 CS generation experiments
With EN-CS as our starting point, we frame CS generation as a machine translation task, with English as the source and CS as the target language, where parts of the source sentence have to be translated to Spanish. In our experiments, we fine-tune four small-sized generative models, namely, Llama3 8B, Llama3 Instruct 8B Dubey et al. (2024), Mistral 7B and Mistral Instruct 7B Jiang et al. (2023).
To fine-tune the models, we use the causal language modelling objective, but with appropriate input formats for the base and instruct models. For base models we use templates (Zhu et al., 2024) in the form of “<X>=<Y>”, where <X> and <Y> are placeholders for the input English sentence and generated CS, respectively. At inference, the second code-switched part is left empty for the model to fill. For instruction-tuned models, we provide a system prompt with the instruction, a query by the user in English, and an answer from the assistant with the code-switched target. At inference time, the answer is left blank (See Table 7 in Appendix D for example prompts).
@@ 4 CS generation experiments
Mistral damm everyone got married y ni somos esa lol ofrever alone
Mistral Instruct dude todo el mundo se casó y nosotros hemos quedado solteros lol
@@
Early experiments indicated that fine-tuned models usually produce the desired output up to a punctuation mark and then either begin to translate the sentence again or hallucinate more content. We therefore truncate the output up to a punctuation mark where the length is closest to that of the original sentence (Bawden and Yvon, 2023).
As baselines, we include few-shot experiments by directly prompting GPT-4o and Llama3.3 70B Instruct to generate CS text using a 5-shot approach (see Table 7 in the Appendix). We refer to these systems in the experiments as GPT-4ofs and Llama3.3-70Bfs, respectively. We also include a strong dedicated MT baseline, developed by fine-tuning the NLLB team et al. (2022) model (nllb-200-distilled-600M) using EN-CS. The model was trained with standard settings.  [^3] A batch size of 32, learning rate of $1\times 10^{-4}$, using constant learning rate schedule with 1,000 warmup steps, a gradient clipping threshold of 1.0, and a weight decay of $1\times 10^{-3}$. Training was conducted for 50,000 steps. For evaluation, we selected the checkpoint that achieved the highest BLEU score on the development set.
5 Qualitative evaluation
5.1 Preference based evaluation
We perform a tournament-based evaluation that allows us to determine the ranking of models in terms of human preference. A total of $880$ instances are matched against each other, corresponding to the outputs of the seven models for $110$ English source sentences, as well as the gold standard reference. The evaluation is conducted pairwise, requiring annotators to choose the best out of two sentences or declare a tie. When choosing the best sentence, annotators do not know the original English sentence, nor which model produced what output. This process results in $080110\cdot{8\choose 2}=3,$ 080 comparisons, and was carried out by $14$ annotators, with each annotator performing at most $300$ random comparisons.
Annotators are provided with a series of criteria to choose between the instances, based on the error analysis described in the next section. They must take into account three main criteria, which must be applied in the following order: a) the presence and naturalness of the CS; b) the content and fluency of the sentences; and c) the orthographical errors of the instances (correct punctuation, presence of typos, etc.). Annotators are furthermore asked to avoid declaring ties, unless completely necessary (e.g., in a case where both sentences are completely monolingual and therefore equally incorrect) to compel them to develop a preference. The complete annotation guidelines are available in Appendix E. Inter-annotator agreement on a subset of 100 sentence pairs shows substantial agreement ($\kappa=0.74$).
Human GPT
Model Score Rank Score Rank
Gold Standard 525.0 [Uncaptioned image] 484.5 [Uncaptioned image]
Llama3 423.0 [Uncaptioned image] 374.0 4
Llama3 Instruct 393.0 [Uncaptioned image] 367.5 5
NLLB 391.0 4 304.0 8
Mistral 368.0 5 334.5 6
GPT-4ofs 366.5 6 478.0 [Uncaptioned image]
Llama3.3-70Bfs 307.0 7 428.0 [Uncaptioned image]
Mistral Instruct 306.5 8 309.5 7
Table 4: Ranking of models according to human preference (c.f. Section 5.1) and using GPT as a judge (c.f. Section 6.2).
We calculate a global score for each model, as follows: every time a model is voted, it gets $1$ point, and the loser gets 00 points; in case of ties, both models get $0.5$ points each. Table 4 shows the global scores, as well as the ranking of human preferences according to said score (second and third columns). We find that the gold standard reference obtains the highest score, as expected, and that fine-tuned Llama3 ranks the highest among the automatic methods. Instruction-tuned models obtain worse scores compared to their base model counterparts, with a similar difference for both Llama3 and Mistral family of models. The NLLB model ranks higher than the Mistral models, but lower than the fine-tuned Llama3 models, showing their potential to rival dedicated models in generation tasks. According to these preferences, fine-tuning LLMs for CS generation can be critical to ensure better results, since the larger models with few-shot prompting rank only above Mistral Instruct, with GPT-4ofs outranking Llama3.3-10Bfs.
5.2 Error analysis
In order to further explore differences between model performance, we analyse the most common errors made by the CS generation models, both quantitatively and qualitatively. We extend the machine translation error typology presented in Popović (2018) to CS generation error analysis. To do so, we randomly select a set of 100 outputs from all models and conduct a detailed examination of the types of errors present in them. This thorough analysis allows us to identify recurring patterns and propose a refined error typology specifically for automatic CS generation. This initial error analysis yields 18 total error categories, which we simplify and group into three main error types: a) CS errors, b) Translation Errors, and c) Format errors. The full error typology, along with detailed descriptions for each error type, is provided in Appendix F, while here we explain the three error categories:
  • CS Errors: Errors of sentences that are either completely monolingual or switch between languages in an unnatural manner, e.g., by repeating the same word in English and Spanish. In Example 5.2, Llama3 Instruct preserves the original meaning, but the sentence is fully monolingual.
@@ 5.2 Error analysis
    Source After all these things when we’re done.
    Output after all these things when we’re finished
[…]
  • Translation errors: Critical errors that either change the original meaning of the sentence or introduce mistakes in fluency or grammar, for example, using the wrong tense or word order. Example 5.2 shows an instance where Mistral Instruct outputs a seemingly natural code-switched sentence, but the phrase “they got hurt” is not adequately translated and the meaning of the sentence is not preserved.
@@
    Output no estuve happy porque me dieron mal
@@
Figure 1: Error distribution by model, obtained by counting the number of instances that present errors of each type.
We classify $700$ additional instances ($100$ instances per model, obtained from the same source sentences) into these kind of errors, and show the results in Figure 1.
GPT-4ofs makes the fewest errors overall (50), closely followed by NLLB (56). However, 90% of GPT’s errors (45) and 45% on NLLB’s (25) are CS related, indicating that while these systems preserve the meaning of sentences and generate few formatting errors, they often produce entirely monolingual outputs, which is a critical error. In comparison, CS-related mistakes are the least common in fine-tuned LLMs, accounting for less than $15\%$ of the overall error count. This analysis show that fine-tuned LLMs have effectively learned to switch between languages naturally, though they may still be prone to other less critical types of errors.
Among the fine-tuned LLMs, Llama models present 19 fewer errors on average than the Mistral family. Base models of both families struggle mainly with format errors, which make up $50.68\%$ of their errors on average, whereas instruction-tuned models present more meaning-related issues, $53,45\%$. This suggests that the linguistic knowledge of the models degrade when tuned on instructions, a phenomenon that has been observed on other related areas Fu et al. (2024).
6 Automatic Evaluation

### v3

@@ Abstract
Code-switching (CS) is still a critical challenge in Natural Language Processing (NLP), due to the limited availability of large-scale, diverse CS datasets for robust training and evaluation. Despite recent advances, the capabilities and limitations of LLMs in handling CS are still not fully understood. In this work, we investigate the extent to which LLMs can be used in a framework for CS text generation, focusing on the English-Spanish language pair. Our proposed methodology consists of back-translating natural CS sentences into monolingual English, and using the resulting parallel corpus to fine-tune LLMs to turn monolingual sentences into CS. We thoroughly analyze the models’ performance through a study on human preferences, a qualitative error analysis, an evaluation with popular reference-based metrics and LLM-based judgment. Results show that fine-tuning can be a key step to ensure that current LLMs consistently generate fluent code-switched text and that our methodology generates high-quality outputs, expanding research opportunities in CS communication. We find that traditional metrics do not correlate with human judgment, and although LLM-based evaluation aligns somewhat more closely, the agreement remains limited. We release our code and generated dataset under a CC-BY-NC-SA license.1
Keywords: Corpus (Creation, Annotation, etc.); Evaluation Methodologies; Multilinguality; Natural Language Generation; Code-switching
[…]
Conditioning LLMs to Generate Code-Switched Text
Maite Heredia   Gorka Labaka   Jeremy Barnes   Aitor Soroa
HiTZ Center - Ixa, University of the Basque Country UPV/EHU
maite.heredia@ehu.eus
Abstract content
11footnotetext: Code [Uncaptioned image]      Dataset[Uncaptioned image]
1.  Introduction
@@
Despite the prevalence of code-switching, most research in Natural Language Processing (NLP) assumes monolingualism as a standard for human communication. However, this implicit decision means that state-of-the-art models are not able to properly interpret or generate CS data. Even advances in multilingual language modelling (Lin et al., 2022; Chowdhery et al., 2023) have not led to significant improvements, and performance on CS data is still poor compared to performance on monolingual data (Aguilar et al., 2020; Winata et al., 2021). This occurs because there is little CS text available in the multilingual pretraining data. Similarly, there are no parallel datasets available to learn to generate CS in a supervised fashion, as one would expect for tasks such as Machine Translation (MT). Finally, existing methodology for evaluating automatically generated CS text, which has specific needs different from other text generation tasks, are still not good enough and fail to capture nuances of CS text (Srivastava and Singh, 2021). It is therefore crucial to develop methodologies to enable models to generate natural CS text and simultaneously implement robust evaluation frameworks that can assess how well NLP systems handle CS across multiple tasks. We argue that both of these goals require models that can conditionally generate CS from monolingual text. Consequently, our research focuses on the development of a methodology to fine-tune and evaluate LLMs on the task of CS generation, following three main research questions:
@@
Do automatic metrics for Natural Language Generation (NLG) or LLM judges correlate well with human judgment for the task of CS generation?
Based on these research questions, we propose a novel approach to generate CS from monolingual text using LLMs and apply it to the English-Spanish pair. We create a new parallel English–CS corpus, EN-CS, by leveraging natural CS data and using LLMs to perform back-translation from CS into English, resulting in high-quality pseudo-parallel pairs, suitable for training and evaluating models on CS generation (RQ2). We provide a comprehensive comparison of CS generation using LLMs in both zero-shot and fine-tuned settings, and we compare their performance against that of a dedicated MT model (RQ1). Finally, we evaluate our methodology both qualitatively, with a study on human preferences and a manual error analysis, and quantitatively, using automatic NLG metrics and LLM as a judge, which allows us to study the correlation between human and automatic evaluation for this task (RQ3). The evaluation is conducted in both in-domain and out-of-domain settings.
2.  Related Work
@@
Most code-switched data stems from social media, while other popular data sources include recordings and transcriptions (Winata et al., 2023). Shared tasks using such CS data have been organized for the tasks of Language Identification (Solorio et al., 2014; Molina et al., 2016) and Sentiment Analysis (Patwa et al., 2020). Similarly, two benchmarks exist to evaluate model performance on CS text, covering different language pairs and tasks: LINCE (Aguilar et al., 2020), which covers tasks such as Part Of Speech tagging or Sentiment Analysis; and GLUECoS (Khanuja et al., 2020), which focuses on NLU tasks for Hindi-English. GLUECoS cannot be currently used without access to the X API.
@@
CS generation has seldom been tackled in previous research. Approaches include linguistically informed techniques to find plausible switching points (Pratapa et al., 2018; Gupta et al., 2020; Gregorius and Okadome, 2022; Hsu et al., 2023; Potter and Yuan, 2024), data augmentation techniques (Tarunesh et al., 2021) and, more recently, prompting LLMs for CS generation (Yong et al., 2023; Terblanche et al., 2024). While CS generation is often evaluated by human annotators (Tarunesh et al., 2021; Gregorius and Okadome, 2022), there remains a need for more robust automatic evaluation methodologies to carefully assess the naturalness and fluency of the generated texts, as recently explored by JudgeLLMs (Kuwanto et al., 2024).
3.  Parallel Data Creation
@@
Train Dev Test Test (ood)
Original 94,728 19,574 33,361 352
Pre-processed 12,933 2,461 5,353 254
EN-CS 10,703 791 1,040 171
@@
3.1.  The LINCE benchmark
[…]
estaba aquí three feet away .
spa spa eng eng eng eng&spa
[…]
LINCE comprises around $95,000$ train, $20,000$ development, and $33,000$ test instances for the English-Spanish pair. We deduplicate the instances among splits, and filter and pre-process the instances to ensure that they are suitable for our task by removing links, replacing usernames with the placeholder <user>, and detokenizing all instances with the script provided as part of the Moses toolkit (Koehn et al., 2007). After this preprocessing, we obtain a more natural version of the LINCE data. A preliminary analysis reveals that many sentences in LINCE are monolingual or contain a single word in one language, which often corresponds to a borrowing, as shown in Example 3.1. We adopt a simple heuristic to approximate the distinction between actual CS and cases that involve only a borrowing or are incorrectly labeled. Specifically, we retain only sentences that contain at least two words in each language, which substantially reduces the likelihood of including cases where an isolated borrowing is tagged as CS.
[…]
After these pre-processing and filtering steps, we end up with $12,933$ train, $2,461$ development and $5,353$ test instances. The comparison between the original size of LINCE and the final number of sentences selected for our experiments after pre-processing is shown in Table 1.
3.2.  EN-CS
The next step in our method requires creating a pseudo-parallel English-CS dataset by translating the natural code-switched instances into monolingual text. As there are no available MT systems to convert from English-Spanish CS text to English monolingual text, we instead make use of prompt engineering, using the Command R model (Cohere For AI, 2024), one of the strongest publicly available models at the time.
We perform an initial set of experiments to determine the optimal prompt to generate monolingual English versions of the code-switched data. Ideally, we aim for a prompt that generates translations that maintain the meaning of the original sentences, are fluent and natural, whose grammar is correct, and that do not contain any Spanish words or phrases. After extensive testing (see Appendix A), we use the following prompt in a 5-shot setting: Now convert this code-switched phrase to English. Leave the parts in English as they are, focus on translating the parts in Spanish. Finally, we filter output instances that contain profanity that was not present in the source texts or irrelevant information, such as Of course, here’s your translation:, because preliminary experiments show that these instances were problematic for the generation task and conditioned the outputs too much.
In order to create a valid gold standard test set, we perform a manual post-edition of the the monolingual test translations for $1,040$ instances of the LINCE test set. The post-edition was carried out by three proficient speakers of English and Spanish, who were provided with specific guidelines, as shown in Appendix B.
@@ 3.2.  EN-CS
3.2.1.  Quality assessment
To estimate the quality of the post-edition process, we compare the post-editions of two annotators on $100$ additional random instances. The results show that $75\%$ of the sentences remain unchanged, as they are already adequate. There is a $87.87\%$ similarity between the post-editions of the two annotators, as measured by Levenshtein distance, demonstrating a high degree of consistency and quality in the post-edition process.
4.  Experimental settings
Generation settings
With EN-CS as our starting point, we frame CS generation as a MT task, with English as the source and CS as the target language, where parts of the source sentence have to be translated to Spanish. In our experiments, we fine-tune two small-sized generative models from the Llama family, namely, Llama3 8B and Llama3 Instruct 8B Dubey et al. (2024).
To fine-tune the models, we use the causal language modelling objective, but with appropriate input formats for the base and instruct models. For the base model, we use templates (Zhu et al., 2024) in the form of “<X>=<Y>”, where <X> and <Y> are placeholders for the input English sentence and generated CS, respectively. At inference, the second code-switched part is left empty for the model to fill. For the instruction-tuned model, we provide a system prompt with the instruction, a query by the user in English, and an answer from the assistant with the code-switched target. At inference time, the answer is left blank (See Appendix D for example prompts).
Models are trained using Quantized Low-Rank Adaptation (QLoRA) (Dettmers et al., 2023) with standard parameters: the model is loaded in 4 bit with NF4 quantization data type and bf16 computational data type. The LoRA rank and scaling factor are set to 16 and the dropout to $0.05$. We apply the LoRA update matrices to the attention blocks and do not train bias parameters. Regarding the hyperparameters, we only tune the learning rate ($1e\textsuperscript{-4}$, $5e\textsuperscript{-4}$, $1e\textsuperscript{-3}$ and $5e\textsuperscript{-3}$) and training epoch $\in[1\ldots 10]$, choosing the parameters that give the lowest cross-entropy loss on the development set for each model. We use the transformers package (Wolf et al., 2020) for all training experiments.
Early experiments indicated that fine-tuned models usually produce the desired output up to a punctuation mark and then either begin to translate the sentence again or hallucinate more content. We therefore truncate the output up to a punctuation mark where the length is closest to that of the original sentence (Bawden and Yvon, 2023). Although simple, this has proven to be the method that yields the best results. We additionally experimented with different generation parameters ( length_penalty and exponential_length_decay), as well as trying to control the length of the generation with length codes, but find that the truncation heuristic performs the best. Accordingly, all further experiments will use the truncated output.
In-domain and out-of-domain evaluation
To evaluate the performance of models, we test them on the test set of EN-CS, and to evaluate their capabilities outside of the domains covered by LINCE, we also propose an out-of-domain evaluation. For this purpose, we gather a series of small creative non-fiction texts originally written using English-Spanish CS (Dwyer, 2017, 2019a, 2019b, 2019c, 2019d, 2020). These are quite different from the other split of the test set in domain and register, as well as more superficial features such as the length of the sentences, that tend to be much longer than those in LINCE. To process this collection of texts in a similar manner to the instances of LINCE, we first divide them into sentences and then obtain the language ID for each token using the models from the CodeSwitch repository. [^3] https://github.com/sagorbrur/codeswitch With each sentence tokenized and tagged with its corresponding language, we can then process and filter the instances just like those sourced from LINCE (c.f. Section 3.1).
Baselines
We include few-shot experiments by directly prompting GPT-4o and Llama3.3 70B Instruct to generate CS text using a 5-shot approach (See Appendix D for example prompts). We refer to these systems in the experiments as GPT-4ofs and Llama3.3-70Bfs, respectively. We also include a strong dedicated MT baseline, developed by fine-tuning the NLLB NLLB Team et al. (2022) model (nllb-200-distilled-600M) using EN-CS. The model was trained with standard settings. [^4] A batch size of $32$, learning rate of $1\times 10^{-4}$, using constant learning rate schedule with 1,0001{,}000 warmup steps, a gradient clipping threshold of $1.0$, and a weight decay of $1\times 10^{-3}$. Training was conducted for 50,000steps50{,}000steps. For evaluation, we selected the checkpoint that achieved the highest BLEU score on the development set.
5.  Qualitative evaluation
5.1.  Preference based evaluation
We perform a tournament-based evaluation that allows us to determine the ranking of models in terms of human preference. A total of $1260$ instances are matched against each other, corresponding to the outputs of the five models for $210$ English source sentences, as well as the gold standard reference. The evaluation is conducted pairwise, requiring annotators to choose the best out of two sentences or declare a tie. When choosing the best sentence, annotators do not know the original English sentence, nor which model produced what output. This process results in $210\cdot{6\choose 2}=3150$ comparisons, and was carried out by $14$ annotators, with each annotator performing at least $100$ random comparisons.
Annotators are provided with a series of criteria to choose between the instances, based on the error analysis described in the next section. They must take into account three main criteria, which must be applied in the following order: a) the presence and naturalness of the CS; b) the content and fluency of the sentences; and c) the orthographical errors of the instances (correct punctuation, presence of typos, etc.). Annotators are furthermore asked to avoid declaring ties, unless completely necessary (e.g., in a case where both sentences are completely monolingual and therefore equally incorrect) to compel them to develop a preference. The complete annotation guidelines are available in Appendix E. Inter-annotator agreement on a subset of $200$ sentence pairs shows substantial agreement for the in-domain set ($\kappa=0.61$) and moderate agreement for the out-of-domain set ($\kappa=0.51$).
In domain Out of domain Total
Model Score Rank Score Rank Score Rank
Gold Standard 369.5 [Uncaptioned image] 434.5 [Uncaptioned image] 804.0 [Uncaptioned image]
Llama3 291.5 [Uncaptioned image] 282.0 [Uncaptioned image] 573.5 [Uncaptioned image]
Llama3 Instruct 270.5 [Uncaptioned image] 210.0 4 480.5 4
NLLB 259.5 4 247.5 [Uncaptioned image] 507 [Uncaptioned image]
GPT-4ofs 251.5 5 162.0 6 413.5 5
Llama3.3-70Bfs 207.5 6 164.0 5 371.5 6
Table 4: Ranking of models according to human preference.
We calculate a global score for each model, as follows: every time a model is voted, it gets $1$ point, and the loser gets 0 points; in case of ties, both models get $0.5$ points each. Table 4 shows the global scores, as well as the ranking of human preferences according to said score (second and third columns). We find that the gold standard reference obtains the highest score, as expected, with an specially stark difference in the out-of-domain set, highlighting the difficulty of this domain for all models. Fine-tuned Llama3 ranks the highest among the automatic methods in both in-domain and out-of-domain settings, which shows its ability to generalize to more challenging domains. The instruction-tuned model Llama3 Instruct obtains worse scores compared to its base model counterpart, which is likely due to instruction tuning reducing certain model capabilities Li et al. (2024); West and Potts (2025). It ranks higher than the NLLB model in-domain, but lower in the out-of-domain evaluation and the overall scores. Still, the difference between Llama3 and NLLB shows the potential to rival dedicated models in generation tasks. According to these preferences, fine-tuning LLMs for CS generation can be critical to ensure better results, since the larger models with few-shot prompting rank, both GPT-4ofs and Llama3.3-70Bfs, are outranked by all models in both evaluation settings.
Figure 1: Error distribution by model, obtained by counting the number of instances that present errors of each type. Analyzed both for the in-domain (In) and out-of-domain (Out) sets.
5.2.  Error analysis
In order to further explore differences between model performance, we analyze the most common errors made by the CS generation models, both quantitatively and qualitatively. We extend the MT error typology presented in Popović2018errorclassification to CS generation error analysis. To do so, we randomly select a set of $100$ outputs from all models and conduct a detailed examination of the types of errors present in them. This thorough analysis allows us to identify recurring patterns and propose a refined error typology specifically for automatic CS generation. This initial error analysis yields $18$ total error categories, which we simplify and group into three main error types: a) CS errors, b) Translation Errors, and c) Format errors. The full error typology, along with detailed descriptions for each error type, is provided in Appendix F.
  • CS Errors: Errors of sentences that are either completely monolingual or switch between languages in an unnatural manner, e.g., by repeating the same word in English and Spanish. In Example 5.2, Llama3.3-70Bfs preserves the original meaning, but the sentence is fully monolingual.
@@ 5.2.  Error analysis
    Source yea… the best i can do is be here for him if he needs me
    Output sí… lo mejor que puedo hacer es estar aquí para él si me necesita
@@
  • Translation errors: Critical errors that either change the original meaning of the sentence or introduce mistakes in fluency or grammar, for example, using the wrong tense or word order. Example 5.2 shows an instance where Llama3 Instruct outputs a seemingly natural code-switched sentence, but the phrase “they got hurt” is not adequately translated and the meaning of the sentence is not preserved.
@@
    Output i wasn’t happy because me dolieron
@@
We classify $1,000$ additional instances ($200$ instances per model, obtained from the same source sentences) into these kind of errors, and show the results in Figure 1.
For the in-domain test set, GPT-4ofs makes the fewest errors overall ($50$), closely followed by NLLB ($56$). However, $90\%$ of GPT’s errors ($45$) and $45\%$ on NLLB’s ($25$) are CS related, indicating that while these systems preserve the meaning of sentences and generate few formatting errors, they often produce entirely monolingual outputs, which is a critical error. In comparison, CS-related mistakes are the least common in fine-tuned LLMs, accounting for less than $15\%$ of the overall error count. This analysis shows that fine-tuned LLMs have effectively learned to switch between languages naturally, though they may still be prone to other less critical types of errors. The Llama3 base model struggles with maintaining the format of the sentences, which makes up $50.68\%$ of its errors, whereas the instruction-tuned model Llama3 Instruct presents more meaning-related issues, $53,45\%$. This suggests that the linguistic knowledge of the model has degraded when tuned on instructions, a phenomenon that has been observed on other related areas Fu et al. (2024).
BLEU BERTScore chrF
Model In domain Out of domain In domain Out of domain In domain Out of domain
Llama3 8B 34.49 14.93 81.64 75.44 53.17 38.26
Llama3 8B Instruct 33.42 14.59 81.77 74.99 52.01 38.62
Llama3.3-70Bfs 22.41 14.10 79.77 78.33 44.57 37.37
GPT-4ofs 32.25 15.65 83.09 78.96 50.48 38.62
NLLB 35.56 14.45 84.11 76.53 54.74 38.38
Identity 33.34 41.54 82.31 83.31 45.51 58.03
When evaluated on a different domain, the Llama3 Instruct and NLLB models produce notably more errors, suggesting that they are overfitting on the train set or have limited ability to generalize to more challenging domains. In particular, NLLB duplicates the number of translation errors. It is also noteworthy that Llama3.3 70Bfs shows a similar increase in errors, despite not being fine-tuned—a behavior that seems to arise due to the domain and longer sentences. This is not the case for Llama3 or GPT-4ofs, which output a similar number of errors for both domains. It is especially interesting that the Llama3 base model generalizes so well, despite extensive fine-tuning. Still, the small size of the out-of-domain test set calls for caution when interpreting these differences.
6.  Automatic Evaluation